# Data-Adaptive Identification of Subpopulations Vulnerable to Chemical Exposures using Stochastic Interventions

### v1

@@ Abstract
In environmental epidemiology, identifying subpopulations vulnerable to chemical exposures and those who may benefit differently from exposure-reducing policies is essential. For instance, sex-specific vulnerabilities, age, and pregnancy are critical factors for policymakers when setting regulatory guidelines. However, current semi-parametric methods for heterogeneous treatment effects are often limited to binary exposures and function as black boxes, lacking clear, interpretable rules for subpopulation-specific policy interventions. This study introduces a novel method using cross-validated targeted minimum loss-based estimation (TMLE) paired with a data-adaptive target parameter strategy to identify subpopulations with the most significant differential impact from simulated policy interventions that reduce exposure. Our approach is assumption-lean, allowing for the integration of machine learning while still yielding valid confidence intervals. We demonstrate the robustness of our methodology through simulations and application to NHANES data. Our analysis of NHANES data for persistent organic pollutants on leukocyte telomere length (LTL) identified age as the maximum effect modifier. Specifically, we found that exposure to 3,3’,4,4’,5-pentachlorobiphenyl (pcnb) consistently had a differential impact on LTL, with a one standard deviation reduction in exposure leading to a more pronounced increase in LTL among younger populations compared to older ones. We offer our method as an open-source software package, EffectXshift, enabling researchers to investigate the effect modification of continuous exposures. The EffectXshift package provides clear and interpretable results, informing targeted public health interventions and policy decisions.
@@ 1 Introduction
Identifying vulnerable subpopulations in environmental health research is crucial for understanding how different groups are differentially impacted by chemical exposures and for informing targeted public health interventions. Numerous studies have highlighted that certain populations, such as children, pregnant women, the elderly, and socioeconomically disadvantaged groups, bear a disproportionate burden of environmental pollutants, leading to significant health disparities Balbus and Malina (2009); Ruiz et al. (2018); Zorzetto et al. (2024a).
For example, research consistently shows that African Americans, Latinos, and low-income individuals are more likely to experience higher exposures to endocrine-disrupting chemicals (EDCs) such as polychlorinated biphenyls (PCBs) and bisphenol A (BPA), which are associated with metabolic disorders, including diabetes Ruiz et al. (2018). These groups are disproportionately exposed due to factors such as dietary patterns, the use of consumer products, and residential proximity to pollution sources, particularly air pollution Makri and Stilianakis (2008). Social determinants of health, such as systemic oppression and reduced access to resources, drive these exposure disparities. Additionally, pregnant women and children are particularly susceptible to the toxic effects of EDCs and exposure to metals, especially lead (Pb) Rolfo et al. (2020); Shah-Kulkarni et al. (2020), which can exacerbate health problems such as neurodevelopmental disorders and cardiovascular disease Healey (2009). The per capita healthcare costs associated with health issues related to EDC exposure are significantly higher in these communities, highlighting the urgent need to address this environmental justice crisis Weiss et al. (2023).
Effect modification is a critical concept in epidemiological research, often assessed to understand how different baseline factors influence the relationship between exposure and outcome. Common methods for assessing effect modification include stratified analysis, interaction terms in regression models, regional GIS estimates, interpreting knot points in spline-based or generalized additive models (GAMs), and Bayesian profile regression. For example, Samoli et al. (2011) examined the acute effects of air pollution on pediatric asthma exacerbation by stratifying the analysis by season, sex, and age, using Poisson regression models with interaction terms to investigate the differential effects of pollutants under various conditions, such as during desert dust days. Similarly, Analitis et al. (2014) explored the modification of the effects of heat waves on mortality by air pollution within the EuroHEAT project, employing generalized estimating equation models to assess the joint effects of high temperatures and air pollution on mortality in multiple European cities.
Bayesian methods have shown significant promise in estimating subpopulation effects within environmental health research, particularly with multipollutant exposures. For example, Coker et al. (2018) demonstrated the efficacy of Bayesian profile regression in identifying subpopulations susceptible to different mixtures of pollutants. Zorzetto et al. (2024b) recently introduced Confounder-Dependent Bayesian Mixture Modeling to characterize heterogeneity of causal effects in air pollution studies, leveraging dependent Dirichlet processes to identify population groups with distinct exposure-response relationships for binary exposures.
Recent advancements in heterogeneous effect estimation have also progressed within the domain of semi-parametric statistics. However, many of these meta-learners, such as the T-learner, the S-learner, and the X-learner, often focus on single binary exposures, which are less applicable to environmental epidemiologists dealing with multiple and continuous exposures Künzel et al. (2019). Despite their theoretical strengths, these methods are limited in environmental studies because of their design focus on binary treatments rather than the complex, multidimensional exposures typically encountered. In addition, many act as black boxes for which the analyst chooses which subgroups of individuals to assess treatment effects for, which is often unknown a priori.
The goal of this paper is to introduce a versatile method for causal inference that achieves two key objectives: (i) identifying the exposure-covariate relationships with the maximal differential impact from an intervention, and (ii) estimating the subpopulation intervention effect in the data-adaptively determined subregion using an assumption-lean estimator. This approach integrates machine learning with data-adaptive targeted parameter estimation to provide robust estimates and valid confidence intervals in covariate regions Hub (2016); Zheng and van der Laan (2010). Our methodology is applicable to any type of exposure, including mixed exposures, by employing stochastic shift interventions Muñoz and van der Laan (2012). These interventions alter the exposure distribution and evaluate potential outcomes under this new distribution, extending to mixtures where the exposure or pairs of exposures with the most significant effect modification are identified and estimated.
By simultaneously identifying these relationships and simulating an intervention, our approach necessitates cross-validation to avoid overfitting to our sample data. In this process, the maximal effect modifier is identified in one part of the data, while the effects of policy intervention within the subregion are estimated on a different part of the data. This ensures that our estimates are both reliable and generalizable. If the maximal effect modifier relationship is consistent across data splits, a pooled estimate is calculated leveraging the full data even though the parameter of interest was discovered from the data.
@@
This paper is structured as follows. Section 2 introduces stochastic shift interventions, providing the foundational concepts and notation. In Section 3, we discuss our oracle target parameter, which defines the maximal effect modifier relationship. Section 4 describes the recursive partitioning algorithm designed to estimate this oracle parameter, identifying subregions within the covariates that exhibit the maximal differential impact of a stochastic intervention. Section 5 presents simulations of a mixed exposure scenario, highlighting the identification and accurate estimation of the maximal effect modifier when applying our method. In Section 6, we apply our approach to NHANES data to explore the effects of mixed POP on telomere length, assessing possible effect modification. The code for our method, along with the applications and simulations, is available on GitHub in our open source package called EffectXshift.
2 Stochastic Shift Interventions
Estimating the causal effect of an intervention on a population typically involves defining parameters in a nonparametric structural equation model Pearl (2000), where the treatment or exposure is assigned deterministically. However, in many real-world scenarios, especially in environmental health, exposures are stochastically assigned due to policy interventions that alter the distribution of exposures rather than assigning fixed values. To account for this, we define our new causal parameter for effect modification based on stochastic interventions.
A stochastic shift intervention modifies the distribution of the exposure rather than setting it to a fixed value. This is particularly relevant in environmental health, where interventions aim to reduce the levels of pollutants, say, reduce all people’s exposure by 2 part per million (ppm) rather than setting exposure to a fixed level, which is not feasible realistically. For example, consider a policy that aims to decrease the levels of mixed POPs (such as DDT, polychlorinated biphenyls (PCBs), and dioxins) in the environment to reduce damage to telomere length. Instead of fixing the exposure to a specific level, the policy shifts the distribution of these POPs downward across the population, for example reduce by 1 standard deviation. Therefore, each person still potentially has a different exposure level, but the entire distribution is shifted by some amount, and we then can compare the expected telomere length under this new exposure distribution to the observed outcome under observed distribution. In our proposed method we assume shifts of $\delta$ are known and fixed already, such as for a particular EPA policy it may be known that a fine reduces emissions by Xon average”, or the analyst is interested in reducing an exposure by 1 unit or 1 standard deviation, etc.
@@ 2 Stochastic Shift Interventions
Consider an experiment where an exposure variable $A$, a continuous or binary outcome $Y$, and a set of covariates $W$ are measured for $n$ randomly sampled subjects. Let $O=(W,A,Y)$ represent a random variable with distribution $P_{0}$, and $O_{1},\ldots,O_{n}$ represent $n$ i.i.d. observations of $O$. The following NPSEM is assumed to hold:
$W=f_{W}(U_{W});\quad A=f_{A}(W,U_{A});\quad Y=f_{Y}(A,W,U_{Y}),$ (1)
where $U_{W}$, $U_{A}$, and $U_{Y}$ are exogenous random variables and $f$ are nonparametric functions. When $U_{A}\perp U_{Y}$, and either $U_{W}\perp U_{Y}$ or $U_{W}\perp U_{A}$ (randomization assumption) is true, the true distribution $P_{0}$ of $O$ can be factorized as:
$P_{0}(O)=P_{0}(Y|A,W)P_{0}(A|W)P_{0}(W),$ (2)
[…]
$P_{\delta}(g_{0})(A=a|W)=g_{0}(a-\delta|W),$ (3)
for a prespecified $\delta$. This represents a shifted version of the current exposure mechanism, where the shift value is constant across strata defined by the covariates.
The objective is to estimate causal parameters under such stochastic interventions. The causal parameter of interest is defined as:
$P(Y_{P_{\delta}}=y)=\int_{a\in\mathcal{A}}\int_{w\in\mathcal{W}}P(Y_{P_{\delta}}=y|A_{P_{\delta}}=a,W=w)\times P_{\delta}(g)(A=a|W=w)P(W=w),da,dw,$ (4)
Where $A_{P_{\delta}}$ denote the exposure variable under the intervened system (i.e., $A_{P_{\delta}}$ is distributed according to $P_{\delta}(g)$) and where $\mathcal{A}$ and $\mathcal{W}$ are the support of A and W, respectively. This is simply the outcome weighted by the new conditional exposure density under a shift.
From the nonparametric structural equation model (NPSEM) given by equation (1), we know that the conditional probability $P(Y_{P_{\delta}}=y|A_{P_{\delta}}=a,W=w)$ is the same as $P(Y_{a}=y|A_{P_{\delta}}=a,W=w)$. Here, $Y_{a}$ represents the counterfactual outcome when the exposure is set to level $a$ with probability one.
Under the usual randomization assumption, which states that $A\perp Y_{a}|W$, we also have that $A_{P_{\delta}}\perp Y_{a}|W$. This implies that $P(Y_{a}=y|A_{P_{\delta}}=a,W=w)=P(Y_{a}=y|W=w)$ and under the consistency assumption (which means that if $A=a$, then $Y_{a}=Y$), the above quantity is identified by $P(Y=y|A=a,W=w)$ our counterfactual distribution can be written as:
$P(Y_{P_{\delta}}=y)=\int_{a\in\mathcal{A}}\int_{w\in\mathcal{W}}P(Y=y|A=a,W=w)\times P_{\delta}(g)(A=a|W=w)\times P(W=w),da,dw,$ (5)
We define the parameter of interest as a mapping $\Psi:\mathcal{M}\rightarrow\mathbb{R}$ that takes an element in a statistical model $\mathcal{M}$ and maps it into a number in the reals. The true value of the parameter is given by the mapping evaluated at the true distribution $P_{0}\in\mathcal{M}$, and is denoted by $\psi_{0}=\Psi(P_{0})$. Our causal and statistical parameter of interest is then given by:
$E(Y_{P_{\delta}})=\Psi(P)=\int_{A\in\mathcal{A}}\int_{W\in\mathcal{W}}\overline{Q}(A,W)\times P_{\delta}(g)(A|W)\times Q_{W}(W),da,dw,$ (6)
To estimate these quantities in practice, we can use ensemble machine learning methods for $g$ and $\overline{Q}$. These advanced techniques allow us to flexibly and more accurately estimate the necessary components of the model.
@@ 2.1 Efficient Estimation and Inference for Stochastic Interventions
To estimate the causal effect of stochastic interventions effectively, we aim to construct an efficient estimator. An efficient estimator is one that achieves the smallest possible asymptotic variance among regular asymptotically linear (RAL) estimators. This ensures that the estimator provides the most precise estimates and the tightest confidence intervals.
An estimator $\hat{\Psi}$ is said to be regular asymptotically linear (RAL) if it satisfies the following property:
@@
where $\phi(O)$ is the influence function. The RAL property implies that the estimator is asymptotically linear, achieving a $\sqrt{n}$ convergence rate, which ensures that the distribution of $\sqrt{n}(\hat{\Psi}-\Psi)$ approaches a normal distribution with mean zero and variance equal to the expectation of the squared influence function. Efficient estimators are those RAL estimators that achieve the smallest possible asymptotic variance among all RAL estimators.
@@
where $\phi(O))$ is the influence function, and $\mathbb{E}_{n}$ denotes the empirical mean. By the central limit theorem, RAL estimators achieve the asymptotic variance bound $\mathbb{E}[\phi^{2}(W)]$. This implies that as the sample size increases, the distribution of $\sqrt{n}(\hat{\Psi}-\Psi)$ approaches a normal distribution with mean zero and variance $\mathbb{E}[\phi^{2}(W)]$, facilitating valid statistical inference.
In the context of our stochastic shift intervention, we seek an estimator $\hat{\Psi}$ for the causal parameter $\Psi$. The discrepancy between the estimator and the true parameter, $\Psi-\hat{\Psi}$, can be characterized using an influence function. The influence function not only quantifies the sensitivity of the estimator to individual observations but also provides a way to derive the asymptotic variance of the estimator, which is crucial for constructing valid confidence intervals.
@@
For our specific stochastic intervention, the influence function $\phi(O)$ is given by:
$\phi(W)=\frac{g(A-\delta|W)}{g(A|W)}\left(Y-\overline{Q}(A,W)\right)+\overline{Q}(A+\delta,W)-\Psi(P).$ (8)
This function represents the deviation of an observation $O=(W,A,Y)$ from the parameter $\Psi(P)$, accounting for the shift $\delta$ in the exposure distribution. Intuitively, the influence function describes how the estimator $\hat{\Psi}$ behaves as we incorporate new observations into our sample. Specifically, the first term $\frac{g(A-\delta|W)}{g(A|W)}\left(Y-\overline{Q}(A,W)\right)$ captures the adjustment for the shift in exposure, weighted by the relative likelihood of the shifted versus original exposure. The second term, $\overline{Q}(A+\delta,W)$, adjusts the expected outcome under the shifted exposure, and the final term $\Psi(P)$ centers the influence function around the true parameter value.
@@
Various estimators can be employed for efficient estimation of our stochastic shift target parameter, including inverse probability weighting (IPW), augmented inverse probability weighting (AIPW), and targeted learning. Among these, we choose targeted learning to estimate our shift parameter due to its robustness in finite samples Li et al. (2022).
The IPW estimator uses the estimated propensity score:
$H(A,W)=\frac{g(A|W)}{g(A-\delta|W)}$ (9)
to weigh observations, allowing for the estimation of causal effects. However, it is inefficient and relies heavily on the correct specification of the propensity score model, which in this case is a density estimator, a particularly difficult estimator where there are not many flexible and reliable methods. As such, this estimator may be unreliable given our need to reliably estimate the conditional density.
The AIPW estimator combines IPW with outcome regression, providing double robustness. It remains consistent if either the propensity score model or the outcome model is correctly specified. However, this estimator is not a plug-in estimator, and therefore does not always respect the bounds of the outcome thereby resulting in potentially more variability.
Targeted Maximum Likelihood Estimation (TMLE) offers a robust and efficient approach by iteratively updating initial estimates through a least favorable submodel. This method adjusts for potential biases in the initial estimates using the clever covariate, ultimately providing more accurate estimates in lower sample sizes. This approach results in a plug-in esitmator which respects the bounds of the data, resulting in lower variability in finite samples.
@@
Targeted learning begins with initial estimates of the outcome regression ($Q$) and the propensity score ($g$), which can be obtained through flexible machine learning techniques. These initial estimates are then updated using a clever covariate, which is constructed to capture the variation related to the shift parameter, $\delta$. The refinement process involves fitting a parametric fluctuation model that minimizes a specified loss function. This iterative process continues until convergence, resulting in a final estimate that is both efficient and less biased. This debiased esitmate is then plugged into the influence function where we can derive Wald-style confidence intervals.

### v2

# Data-Adaptive Identification of Effect Modifiers through Stochastic Shift Interventions and Cross-Validated Targeted Learning

@@ Abstract
In epidemiology, identifying subpopulations that are particularly vulnerable to exposures and those who may benefit differently from exposure-reducing interventions is essential. Factors such as age, gender-specific vulnerabilities, and physiological states such as pregnancy are critical for policymakers when setting regulatory guidelines. However, current semi-parametric methods for estimating heterogeneous treatment effects are often limited to binary exposures and can function as black boxes, lacking clear, interpretable rules for subpopulation-specific policy interventions. This study introduces a novel method that uses cross-validated targeted minimum loss-based estimation (TMLE) paired with a data-adaptive target parameter strategy to identify subpopulations with the most significant differential impact of simulated policy interventions that reduce exposure. Our approach is assumption-lean, allowing for the integration of machine learning while still yielding valid confidence intervals. We demonstrate the robustness of our methodology through simulations and application to data from the National Health and Nutrition Examination Survey. Our analysis of NHANES data on persistent organic pollutants (POPs) and leukocyte telomere length (LTL) identified age as a significant effect modifier. Specifically, we found that exposure to 3,3’,4,4’,5-pentachlorobiphenyl (PCNB) consistently had a differential impact on LTL, with a one-standard deviation reduction in exposure leading to a more pronounced increase in LTL among younger populations compared to older ones. We offer our method as an open-source software package, EffectXshift, enabling researchers to investigate the effect modification of continuous exposures. The EffectXshift package provides clear and interpretable results, informing targeted public health interventions and policy decisions.
@@ 1 Introduction
Environmental risk factors—including pollutants and social determinants—pose significant health risks, but their impact is not uniformly distributed across populations. Certain subpopulations—such as children, pregnant women, the elderly, and socioeconomically disadvantaged communities—are disproportionately exposed to harmful conditions, leading to pronounced health disparities Balbus and Malina (2009); Ruiz et al. (2018); Zorzetto et al. (2024a) For instance, studies have shown that African Americans, Latinos, and low-income individuals are more likely to experience higher exposures to environmental pollutants like air pollution and endocrine-disrupting compounds (EDCs) such as polychlorinated biphenyls (PCBs) and bisphenol A (BPA), as well as adverse social conditions, all of which are linked to metabolic disorders and diabetes Ruiz et al. (2018). Factors contributing to these disparities include dietary patterns, use of certain consumer products, residential proximity to pollution sources, and social determinants like socioeconomic status and access to healthcare Makri and Stilianakis (2008).
Identifying these vulnerable subpopulations is crucial for designing targeted public health interventions and policies aimed at reducing exposure and mitigating health risks. Effect modification—the phenomenon where the effect of an exposure on an outcome differs across levels of another variable - is a key concept in epidemiological research for uncovering such differential impacts Rothman et al. (2008). Traditional methods for assessing effect modification, such as stratified analyses, interaction terms in regression models, and generalized additive models (GAMs), have been widely used Samoli et al. (2011); Analitis et al. (2014). For example, Samoli et al. Samoli et al. (2011) examined how air pollution affects pediatric asthma exacerbations by stratifying analyses by age, sex, and season. However, these conventional methods have limitations when dealing with complex environmental exposures, which are often continuous, multidimensional, and may act synergistically as mixtures Gibson et al. (2019).
While Bayesian methods such as Bayesian profile regression and confounder-dependent mixture modeling have shown promise in identifying subpopulations susceptible to multipollutant exposures Coker et al. (2018); Zorzetto et al. (2024b), they can be computationally intensive for very large datasets and may require specialized expertise to implement and interpret. Similarly, meta-learning methods for estimating heterogeneous treatment effects, like the T-learner, S-learner, and X-learner Künzel et al. (2019), while powerful, often focus on binary treatments and may lack the interpretability desired in environmental studies dealing with continuous exposures. Our approach aims to address these challenges by providing a computationally efficient, interpretable method that naturally handles continuous exposures and leverages machine learning techniques within a causal inference framework.
To address these challenges, we propose a novel, data-adaptive method that leverages causal inference techniques with stochastic shift interventions to identify subpopulations most vulnerable to harmful exposures. Stochastic shift interventions modify the distribution of exposures rather than setting them to fixed values, making them particularly relevant for environmental policies aiming to reduce pollutant levels across populations Muñoz and van der Laan (2012); Kennedy (2019). By integrating targeted maximum likelihood estimation (TMLE) with machine learning, our approach allows for flexible modeling of complex exposure-outcome relationships while providing valid statistical inference van der Laan and Rose (2011).
To illustrate the practical significance of our method, consider the exposure to per- and polyfluoroalkyl substances (PFAS), a group of pollutants found in various consumer products and associated with adverse health outcomes such as thyroid cancer Jain (2015). Suppose that we have biomarker data on PFAS levels from blood samples, thyroid cancer outcomes, and baseline covariates such as age, body mass index (BMI), use of consumer products, and other relevant factors. Our study aims to inform analysts and policymakers about which pollutants in the PFAS mixture has the largest differential impact if reduced in certain subpopulations. For example, we might find that reducing exposure to perfluorononanoic acid (PFNA) by 5 parts per billion (ppb) leads to a 12% decrease in thyroid-related diseases among pregnant women aged 18–29, compared to a 3% decrease in other individuals. Such insights enable the development of targeted interventions and policies that focus on the most vulnerable groups and the most impactful exposures.
Our method achieves two key objectives: (i) identifying the exposure-covariate relationships where an intervention would have the maximal differential impact, and (ii) estimating the effects of such interventions within the identified subpopulations using an assumption-lean estimator. This approach enhances interpretability by providing clear rules for subpopulation-specific interventions and addresses limitations of existing methods that function as black boxes.
@@
This paper is structured as follows. Section 2 introduces stochastic shift interventions, providing the foundational concepts and notation. In Section 3, we discuss our oracle target parameter, which defines the maximal effect modifier relationship. Section 3.3 describes the recursive partitioning algorithm designed to estimate this oracle parameter, identifying subregions within the covariates that exhibit the maximal differential impact of a stochastic intervention. Section 4 presents simulations of a mixed exposure scenario, and 5 shows the identification and accurate estimation of the maximal effect modifier when applying our method to simulatino data. In Section 6.1, we apply our approach to NHANES data to explore the effects of mixed POPs on telomere length, assessing possible effect modification. We close with a conclusion in 7. The code for our method, along with the applications and simulations, is available on GitHub in our open-source package called EffectXshift.
2 Defining Stochastic Shift Interventions
@@
Consider an experiment where an exposure variable $A\in\mathcal{A}$, an outcome $Y$, and a set of covariates $W\in\mathcal{W}$ are measured for $n$ randomly sampled units, with $O=(W,A,Y)$ denoting a random variable drawn from a distribution $P_{0}$. Let $O_{1},\ldots,O_{n}$ be $n$ i.i.d. observations of $O$.
We posit a Nonparametric Structural Equation Model (NPSEM) that encodes the data-generating process:
$W=f_{W}(U_{W}),\quad A=f_{A}(W,U_{A}),\quad Y=f_{Y}(W,A,U_{Y}),$
where $U_{W},U_{A},U_{Y}$ are exogenous random variables and $f_{W},f_{A},f_{Y}$ are deterministic functions. Under this model, the observed data distribution factors as:
$P_{0}(O)=P_{0,W}(W)\,P_{0,A}(A\mid W)\,P_{0,Y}(Y\mid A,W).$
Define a function
$g_{0}:\mathcal{A}\times\mathcal{W}\to\mathbb{R}^{+},$
and use $g_{0}(a\mid w)$ as shorthand for the conditional density of $A$ given $W=w$. Let $\overline{Q}_{0}(a,w)=E_{0}[Y\mid A=a,W=w]$ and let $p_{0,W}(w)$ denote the density of $W$.
We now consider a stochastic shift intervention that modifies the exposure mechanism. Instead of drawing $A$ from $g_{0}(a\mid w)$ as originally defined, we define a shifted exposure distribution:
$p_{0,A_{\delta}}(a\mid w)=g_{0}(a-\delta\mid w),$
where $\delta\in\mathbb{R}$ characterizes the magnitude and direction of the shift. Under this modified NPSEM, each unit’s exposure is effectively shifted by $\delta$ units, and the corresponding counterfactual outcome $Y_{A_{\delta}}$ is the outcome that would have been observed if the unit’s exposure had been drawn from the shifted distribution $p_{0,A_{\delta}}$.
Our parameter of interest is:
$\Psi(P_{0})=E_{P_{0,A_{\delta}}}[Y_{A_{\delta}}],$
the expected outcome under the stochastic shift intervention.
Identification of $\Psi(P_{0})$ from observed data relies on standard causal conditions:
Exchangeability (No Unmeasured Confounding)
Under exchangeability, $Y_{a}\perp A\mid W$ for all $a\in\mathcal{A}$. Given $W$, the observed assignment of $A$ can be viewed as if it were randomized. This assumption ensures that the conditional associations we observe can be interpreted as reflecting causal effects of the exposure on the outcome.
Positivity
The positivity condition requires that the shift $\delta$ is chosen so that the shifted exposure levels remain within the support of $A$ for each covariate stratum. Let $l(w)$ and $u(w)$ be the minimum and maximum values of $A$ attainable for units with $W=w$. To maintain positivity, we must ensure that for all $(a,w)$ in the support of $(A,W)$, the point $a-\delta$ also lies within the interval $[l(w),u(w)]$. If $a-\delta$ falls outside this support, then $g_{0}(a-\delta\mid w)=0$ for those values, violating positivity and preventing identification of $\Psi(P_{0})$.
In practice, this means selecting $\delta$ so that the shifted exposure remains in the feasible range defined by $l(w)$ and $u(w)$. When this is satisfied, there exists an $\epsilon>0$ such that
$0<\frac{g_{0}(a-\delta\mid w)}{g_{0}(a\mid w)}\leq\frac{1}{\epsilon}$
for all relevant $(a,w)$, ensuring that the stochastic shift intervention is well-defined and that the resulting counterfactual distribution $P_{0,A_{\delta}}$ is identifiable from the observed data.
Interpretation
Intuitively, $\Psi(P_{0})$ represents the expected outcome if we replaced the original exposure distribution $g_{0}(a\mid w)$ with one shifted by $\delta$. This parameter captures the causal effect of a hypothetical population-level intervention that systematically reduces exposure levels. By estimating $\Psi(P_{0})$, we can assess the global impact of such an intervention and compare effects across subpopulations defined by $W$, thereby guiding targeted interventions and informing policy decisions. Because the estimation relies only on observational data and these well-defined causal assumptions (exchangeability and positivity), it allows us to infer the potential effects of interventions that have not been experimentally evaluated in practice.
@@ 2.1 Efficient Estimation and Inference for Stochastic Interventions
To estimate the causal effect of stochastic interventions effectively, we aim to construct a semiparametric efficient estimator. Such an estimator achieves the smallest possible asymptotic variance among regular asymptotically linear (RAL) estimators, satisfying:
@@
where $\phi(O)$ is the efficient influence function, and $\mathbb{E}_{n}$ denotes the empirical mean. By the Central Limit Theorem, RAL estimators achieve the asymptotic variance bound $\mathbb{E}[\phi^{2}(O)]$, facilitating valid statistical inference.
@@
In the context of our stochastic shift intervention, the efficient influence function $\phi_{P}(O)$ of $\Psi$ at distribution $P$ is given by:
$\phi_{P}(O)=\frac{g_{0}(A-\delta\mid W)}{g_{0}(A\mid W)}\left(Y-\overline{Q}_{0}(A,W)\right)+\overline{Q}_{0}(A-\delta,W)-\Psi(P).$ (2)
@@
Several estimators can be employed for estimating our stochastic shift parameter, including inverse probability weighting (IPW), augmented inverse probability weighting (AIPW), and targeted maximum likelihood estimation (TMLE). However, IPW estimators can be inefficient and sensitive to model misspecification, and AIPW estimators may not respect the bounds of the outcome, potentially resulting in greater variability. Therefore, we choose TMLE to estimate our parameter due to its efficiency and robustness in finite samples Li et al. (2022).
@@
Targeted learning begins with initial estimates of the outcome regression $\hat{\overline{Q}}(A,W)$ and the exposure density $\hat{g}(A\mid W)$, which can be obtained through flexible machine learning techniques. These initial estimates are then updated using a clever covariate based on the efficient influence function. The refinement process involves fitting a parametric fluctuation model that targets the parameter of interest. This results in a final estimate $\hat{\overline{Q}}^{\star}(A-\delta,W)$ that is efficient and less biased. The updated estimate is then used to compute the TMLE of $\Psi$, and the influence function is utilized to derive confidence intervals.